OpenStudy (nudeoua):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 3 over 4 power, all over 2 to the 1 over 2 power the eighth root of 2 to the third power the square root of 2 to the 3 over 4 power the fourth root of 2 the square root of 2

OpenStudy (zzr0ck3r):

\[\frac{2^{\frac{3}{4}}}{2^\frac{1}{2}}\]?

OpenStudy (zzr0ck3r):

hint \(\dfrac{a^b}{a^c}=a^{b-c}\)

Directrix (directrix):

@nudeoua Do you agree?